SUB-DOPPLER SPECTROSCOPY OF ND_{3}H^{+} ION IN THE NH STRETCHING MODE

\underline{CHIH-HSUAN CHANG}, GRANT T. BUCKINGHAM , and DAVID J. NESBITT, JILA, National Institute of Standards and Technology University of Colorado, and Department of Chemistry and Biochemistry University of Colorado at Boulder, Colorado 80309.

Despite early successes with velocity modulation spectroscopy^{a}^{,}^{b}of ammonium (NH_{4}^{+} ) and perdeuteroammonium (ND_{4}^{+}) cation, corresponding studies of any of the mixed H/D isotopomers (NH_{m}D_{4−m}^{+}) have proven elusive. In this talk, we present first high resolution results on the lone NH stretch fundamental mode for the jet cooled ND_{3}H^{+} ion, based on tunable IR difference frequency absorption spectroscopy in a slit jet, sub-Doppler resolution infrared spectrometer. Supersonically cooled ND_{3}H^{+} ions are generated by modulated (50 KHz) discharges (650V, 500 mA) in ND_{3} doped H_{2}/Ne slit jet expansions, monitored by time-gated, lock-in detection methods and with absorption sensitivities near the quantum shot noise limit. Jet cooled (30 K) P, Q, and R branch rovibrational progressions in the a-type NH stretch band are observed and unambiguously assigned by four line ground state combination differences, with additional confirmation by nuclear spin statistical weights for the 3 identical D atoms (I = 1). Least squares fits to this parallel band yield precision rotational constants and an NH stretch vibrational band origin at 3316.8347(19) cm^{−1}. These high resolution spectroscopic results prove to be in generally excellent agreement with high level ab initio theoretical predictions by Martin and Lee^{c}.

aMark.W. Crofton, Takeshi. Oka, J. Chem. Phys. 86, 5983, (1987).

bE. Schafer, M. H. Begemann, C. S. Gudeman, R. J. Saykally, J. Chem. Phys. 79, 3159, (1983).

cJan. M. L. Martin, Timothy J. Lee, Chem. Phys. Lett. 258, 129 (1996)